# Positive and Unlabeled Data: Model, Estimation, Inference, and Classification

### v2

@@ 5.1 Ablation Study
    1000 11.0% 100% 100% 100% 100% 100% 100% 99.8%
    2000 9.2% 100% 100% 100% 100% 100% 100% 100%
    3000 7.8% 100% 100% 100% 100% 100% 100% 100%
    4000 5.4% 100% 100% 100% 100% 100% 100% 100%
    5000 5.4% 100% 100% 100% 100% 100% 100% 100%
    When $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}_{3}={\bf 0}_{15}$ (i.e., the null hypothesis is true; see column “0”), the rejection rate (i.e., the type-I error) converges to the nominal significant value $0.05$ as sample size $n$ increases, which honors the asymptotic results in Section 4.1. For columns “1” through “7”, the null hypothesis does not hold because $\boldsymbol{\mu}_{3}\neq\boldsymbol{\mu}_{1}$, and the rejection rates (i.e., power) are nearly uniformly 100%, indicating that the proposed hypothesis test almost always rejects the null, even if $\boldsymbol{\mu}_{3}$ and $\boldsymbol{\mu}_{1}$ differ in only one dimension.
@@
    We used two different data-generating mechanisms in this task. The first one generates SCAR PU data, where the source positive (i.e., $p_{s}({\bf x}|y=1)$) and target positive (i.e., $p_{t}({\bf x}|y=1)$) distributions are the same and are set as $\mathrm{MVN}({\bf 0}_{15},\mathbf{I})$; whereas the target negative distribution is set as $p_{t}({\bf x}|y=0)\sim\mathrm{MVN}({\bf 1}_{15},\mathbf{I})$. The second one generates SAR PU data, where the source positive distribution is set as $\mathrm{MVN}({\bf 0}_{15},\mathbf{I})$, the target positive distribution is set as $\mathrm{MVN}(({\bf 1}^{\rm T}_{7}~{}{\bf 0}_{8}^{\rm T})^{\rm T},\mathbf{I})$, and the target negative distribution is $\mathrm{MVN}({\bf 1}_{15},\mathbf{I})$.
@@
    Similar to the previous task, we let $n=m$ and varied $n$ in $\{1000,2000,3000,4000,5000\}$. We experimented on two values of $\pi\in\{0.3,0.7\}$ and repeated each factor level combination for 500 times. We estimated the mixture proportion $\pi=\Pr_{t}(Y=1)$ and constructed confidence intervals for every replicate of each factor level combination using both the DETM and SETM.
@@
    The estimation results are summarized in Table 2 and Figure 2. In the case of the SCAR PU data, both the DETM and SETM provide good estimates. In the left panel of Table 2, we see that the estimated bias, standard deviation, and MSE are all small. Regarding the estimation efficiency, illustrated in the first row of Figure 2, we observe that the SETM outperforms DETM, albeit marginally. The efficiency gain is due to that the SETM is a reduced model of DETM. For the SAR data, the DETM method still performed well with MSE shrinking toward 0 as the sample size increases. However, the SETM failed due to model misspecification as illustrated in the second row of Figure 2 and the right panel of Table 2.
    Figure 2: MSE of ${\pi}$ estimates.
    Table 2: The estimated means and standard deviations (in parenthesis) of $\pi$ estimates.
    $n$ SCAR Data SAR Data SETM DETM SETM DETM $\pi=0.3$ $\pi=0.7$ $\pi=0.3$ $\pi=0.7$ $\pi=0.3$ $\pi=0.7$ $\pi=0.3$ $\pi=0.7$ 1000 0.301 (0.018) 0.702 (0.020) 0.302 (0.020) 0.700 (0.023) 0.001 (0.002) 0.999 (0.002) 0.307 (0.061) 0.669 (0.064) 2000 0.300 (0.012) 0.699 (0.014) 0.300 (0.012) 0.700 (0.015) 0.001 (0.001) 0.999 (0.001) 0.304 (0.036) 0.688 (0.047) 3000 0.300 (0.009) 0.700 (0.011) 0.300 (0.010) 0.700 (0.012) 0.001 (0.001) 1.000 (0.001) 0.304 (0.032) 0.697 (0.040) 4000 0.300 (0.008) 0.699 (0.009) 0.300 (0.009) 0.699 (0.010) 0.001 (0.001) 1.000 (0.001) 0.302 (0.027) 0.694 (0.033) 5000 0.300 (0.007) 0.699 (0.008) 0.300 (0.008) 0.699 (0.008) 0.001 (0.001) 1.000 (0.001) 0.302 (0.023) 0.694 (0.029)
@@
    We also constructed two-sided $95\%$ confidence intervals for $\pi$, and the results are given in Figure 3. We report the estimated coverage probabilities in each of the different scenarios: 1) SETM on the SCAR PU data, 2) DETM on the SCAR PU data, 3) SETM on the SAR PU data, and 4) DETM on the SAR PU data. For the SCAR data, the coverage probabilities of both DETM and SETM reach the nominal level of 95% as the sample size increases. However, when the data are SAR, the confidence intervals constructed using the SETM fail to cover the true parameter value $\pi$, resulting in a coverage probability that is close to 0. On the other hand, the DETM’s confidence intervals still perform well, remaining around the nominal levels of 95%.
    Figure 3: Coverage probabilities of confidence intervals for $\pi$.
@@
    Lastly, we aim to assess the prediction performance of our proposed method. We generate the SCAR and SAR data the same way as in the “Estimations and Confidence Intervals for the Mixture Proportion” part, where $\pi\in\{0.3,0.7\}$ and $n=m\in\{1000,2000,3000,4000,5000\}$. To avoid overfitting, we generate a new validation dataset from the target distribution $p_{t}({\bf x},y)$ with sample size $n$, and the classifiers are applied to the validation dataset to calculate the classification accuracy.
    We compare the following methods, including ours and some existing methods in the machine learning literature. 1) DETM: the proposed approximated Bayes classifier by replacing parameters with the MELEs under the DETM. 2) SETM: the approximated Bayes classifier by replacing parameters with the MELEs under the SETM. 3) PS: proposed by Bekker et al., (2019) and is based on the propensity score function. 4) KM2: proposed by Ramaswamy et al., (2016) using the reproducing kernel Hilbert space and is based on SCAR PU data. 5) TICE: proposed by Bekker and Davis, (2018) and is based on SCAR PU data.
    Figure 4 illustrates the classification accuracies using 100 replicates. When the data are SCAR and $\pi=0.3$, all classifiers performed well, achieving an accuracy exceeding 95% across all sample sizes. However, when $\pi=0.7$, the KM2 method had the lowest accuracy at around 87%, while all other methods maintained at least 90% accuracy across all sample sizes.
@@
    For the SAR data, the simulation studies highlight discrepancies among different methods. In both $\pi=0.3$ and $\pi=0.7$ scenarios, our proposed DETM method achieves the highest classification rate of around 90%, with accuracy improving as the sample size increases. In comparison, at $\pi=0.3$, all benchmark methods have accuracies between 70% and 80%, with SETM performing the worst across all sample sizes. When $\pi=0.7$, the accuracies of benchmark methods PS, KM2, and TICE are around 60%. SETM again has the worst performance.
@@
    The ablation study demonstrates that our proposed DETM method excels in estimation, inference, and classification tasks, regardless of whether SAR or SCAR data are used. Under correct model assumptions, the SETM method performs well, with slightly improved estimation efficiency compared to DETM. However, SETM lacks robustness and performs poorly across all tasks when the SCAR assumption is violated. Additionally, the goodness-of-fit test effectively distinguishes SAR data by accurately identifying SCAR data with a type-I error rate close to the nominal level and exhibiting high power in detecting non-SCAR data. For classification tasks, the proposed DETM method also shows superior performance compared to existing methods.
@@ 5.2 Mobile Phone Dataset
    Before the analysis, we first preprocess the dataset to fit our needs. We first merge classes 0 and 1 to form the positive target data from $p_{t}({\bf x}|y=1)$ (low-end phone). We then use class 2 as the positive source data from $p_{s}({\bf x}|y=1)$ (low-end phone) and class 3 as negative target data from $p_{t}({\bf x}|y=0)$ (high-end phone). Then, the source positive data have a sample size of $n=500$, and the target data have a sample size of $m=1500$, where $500$ are positive and $1000$ are negative. Our formulation ensures that the low-end phones in the source and target are different, as low-end phones in the target data are less expensive; thus, the data can be seen as SAR PU.
    We first conducted the goodness-of-fit test for SETM vs. DETM. The value of the ELR statistic is $R_{N}=1218.628$, which gives a p-value of 0, calibrated by the limiting distribution $\chi^{2}_{20}$. Hence, we reject the null hypothesis with overwhelming evidence and adopt the DETM. The test results also align with the way we construct the PU data. Next, we examine the estimation and inference of the mixture proportion $\pi$. The true $\pi$ in the target data is $\pi_{0}=n_{1}/(n_{1}+n_{3})=2/3$, and the estimate $\widehat{\pi}=0.667$ is close to the true value. For the inference of $\pi$, we examine the behavior of the ELR function $R_{N}^{\ast}(\pi)$ in Figure 5. The vertical solid blue line is the true value of $\pi$. The red horizontal line is the 95% quantile of $\chi^{2}_{1}$, and the two dashed lines indicate the two endpoints of a $95\%$ confidence interval. We can see that the true value is within the confidence interval $[0.6425,0.6903]$.
@@
    Lastly, we focus on the classification. Letting the sample size of the positive target, positive source, and negative target data be $n_{1}=1000,n_{2}=500$, and $n_{3}=500$, respectively, we reserve 20% of the positive target and negative target data and combine them for verification to avoid overfitting. Thus, the PU data have a sample size of $N=n+m$, where $n=n_{2}=500$ and $m=0.8n_{1}+0.8n_{3}=1200$. The sample size of the verification dataset is $0.2n_{1}+0.2n_{3}$. We repeated the data partition on the target data for 20 times and performed classification for each replicate. We compared the proposed DETM method with the benchmarks used in the ablation study. Table 3 summarizes the average classification accuracy from $20$ sample splitting replicates. The numbers in the parenthesis are the standard deviation of the 20 accuracies from 20 replicates. Our proposed method, DETM, has surprisingly achieved perfect classification, whereas the benchmark methods have accuracies around 89.15% to 91.23%.
@@
    Method DETM PS KM2 TICE Accuracy 100% (0%) 91.23% (1.2%) 89.15% (3.23%) 88.42% (1.78%)
@@ 6 Concluding Remarks
    We conclude by highlighting several promising directions for future research. One intriguing extension involves exploring the open set label shift setting (e.g., Garg et al., 2022), where the target domain introduces a new class not present in the source domain, but the source domain has multiple classes. This setting expands the scope of PU data and presents new challenges and opportunities for DETM to demonstrate its adaptability and effectiveness. Another critical area for future investigation is the robustness of DETM in the presence of noisy labels. Enhancing the model’s resilience to such imperfections is vital for ensuring reliable performance in real-world applications, where data quality can often be compromised.
    Appendix: Form of ${\boldsymbol{\Sigma}}$ and Regularity Conditions
    We note that the log-EL function in (12) can be rewritten as $\ell_{N}({\boldsymbol{\theta}})=\min_{\lambda_{1},\lambda_{2}}h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})$, where
@@
    $h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})$ $=$ $-\sum_{i=1}^{N}{\rm log}\left[1+\sum_{t=1}^{2}\lambda_{t}\left\{\exp(\alpha_{t}+{\bf x}_{i}^{\rm T}{\boldsymbol{\beta}}_{t})-1\right\}\right]$ (14)
    Equivalently, $\ell_{N}({\boldsymbol{\theta}})=h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})$ with $(\lambda_{1},\lambda_{2})$ being the solution to $\partial h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})/\partial\lambda_{1}=0$ and $\partial h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})/\partial\lambda_{2}=0.$
    The form of $\Sigma$ depends on the first and second derivatives of $h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})$ with respect to $\lambda_{1}$, $\lambda_{2}$, and ${\boldsymbol{\theta}}$. Specifically, let $\boldsymbol{\upsilon}=(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}}^{\rm T})^{\rm T}$, $\boldsymbol{\upsilon}^{o}=(c\pi^{o},c(1-\pi^{o}),{\boldsymbol{\theta}}^{o\rm T})^{\rm T}$, and $h(\boldsymbol{\upsilon})=h(\lambda_{1},\lambda_{2},{\boldsymbol{\theta}})$. Further, let
    ${\bf V}=N^{-1}\mathbb{E}\left\{\frac{\partial h(\boldsymbol{\upsilon}^{o})}{\partial\boldsymbol{\upsilon}}\frac{\partial h(\boldsymbol{\upsilon}^{o})}{\partial\boldsymbol{\upsilon}^{\rm T}}\right\}\quad\mbox{and}\quad{\bf W}=-N^{-1}\mathbb{E}\left\{\frac{\partial^{2}h(\boldsymbol{\upsilon}^{o})}{\partial\boldsymbol{\upsilon}\partial\boldsymbol{\upsilon}^{\rm T}}\right\}.$ (15)
    It can be verified that both ${\bf V}$ and ${\bf W}$ are independent of the total sample size $N$. The detailed forms of ${\bf V}$ and ${\bf W}$ are given in the supplementary materials. The matrix ${\boldsymbol{\Sigma}}$ is then defined as
    ${\boldsymbol{\Sigma}}={\bf W}^{-1}{\bf V}{\bf W}^{-1}.$ (16)
    The asymptotic results in Section 4 rely on the following regularity conditions. Let ${\bf q}({\bf x})=(1,{\bf x}^{\rm T})^{\rm T}$ and $\mathbb{E}_{s}$ denote the expectation with respect to $p_{s}({\bf x}|y=1)$.
    • C1.
      The total sample size $N=n+m\to\infty$ and $m/N\to c$ for some constant $c\in(0,1)$. Furthermore, $10<\pi^{o}<$ 1, and ${\boldsymbol{\theta}}^{o}$ is an interior point of the parameter space for ${\boldsymbol{\theta}}$.
    • C2.
      The function $\mathbb{E}_{s}\left\{\exp({\bf X}^{\rm T}{\boldsymbol{\beta}})\right\}$ is finite when ${\boldsymbol{\beta}}$ is in a small neighborhood of ${\boldsymbol{\beta}}_{1}^{o}$ and ${\boldsymbol{\beta}}_{2}^{o}$.
    • C3.
      $\mathbb{E}_{s}\left\{{\bf q}({\bf X}){\bf q}({\bf X})^{\rm T}\right\}$ exists and is positive definite.
    • C4.
      ${\bf W}$ has the full rank.
    As we commented before, Condition C1 indicates that both $n$ and $m$ go to infinity at the same rate. Condition C2 guarantees the existence of finite moments of ${\bf q}({\bf X})$ under $p_{s}({\bf x}|y=1)$, $p_{t}({\bf x}|y=0)$, and $p_{t}({\bf x}|y=1)$. Condition C3 is an identifiability condition, and it ensures that the components of ${\bf q}({\bf x})$ are linearly independent under $p_{s}({\bf x}|y=1)$, $p_{t}({\bf x}|y=0)$, and $p_{t}({\bf x}|y=1)$. Hence, the elements of ${\bf q}({\bf x})$ except the first cannot be constant functions. Conditions C1–C4 guarantee that quadratic approximations of the profile log-EL function $\ell_{N}({\boldsymbol{\theta}})$ are applicable.

### v3

@@ 5.1 Simulation Study
1000 10.7% 100% 100% 100% 100% 100% 100% 100%
2000 7.4% 100% 100% 100% 100% 100% 100% 100%
3000 5.3% 100% 100% 100% 100% 100% 100% 100%
4000 5.6% 100% 100% 100% 100% 100% 100% 100%
5000 5.0% 100% 100% 100% 100% 100% 100% 100%
When $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}_{3}={\bf 0}_{15}$ (i.e., the null hypothesis is true; see column “0”), the rejection rate (i.e., the type-I error) converges to the nominal significant value $0.05$ as sample size $n$ increases, which honors the asymptotic results in Section 4.1. For columns “1” through “7”, the null hypothesis does not hold because $\boldsymbol{\mu}_{3}\neq\boldsymbol{\mu}_{1}$. To our surprise, the rejection rates (i.e., power) are uniformly 100%, indicating that the proposed hypothesis test will always reject the null, even if $\boldsymbol{\mu}_{3}$ and $\boldsymbol{\mu}_{1}$ differs in only one dimension.
@@
We use two different data-generating mechanisms in this task. The first one generates CD PU data, where the source positive (i.e., $p_{s}({\bf x}|y=1)$) and target positive (i.e., $p_{t}({\bf x}|y=1)$) distributions are the same and are set as $\mathrm{MVN}({\bf 0}_{15},\mathbf{I})$; whereas the target negative distribution is set as $p_{t}({\bf x}|y=0)\sim\mathrm{MVN}({\bf 1}_{15},\mathbf{I})$. The second one generates DD PU data, where the source positive distribution is set as $\mathrm{MVN}({\bf 0}_{15},\mathbf{I})$, the target positive distribution is set as $\mathrm{MVN}(({\bf 1}^{\rm T}_{7}~{}{\bf 0}_{8}^{\rm T})^{\rm T},\mathbf{I})$, and the target negative distribution is $\mathrm{MVN}({\bf 1}_{15},\mathbf{I})$.
@@
Similar to the previous task, we let $n=m$ and vary $n$ in $\{1000,2000,3000,4000,5000\}$. We use two values of $\pi\in\{0.3,0.7\}$ and repeat each factor level combination for 1000 times. We estimate the mixture proportion $\pi=\Pr_{t}(Y=1)$ and construct confidence intervals using both the DETM and SETM.
@@
The estimation results are summarized in Table 2. In the case of the CD PU data, both the DETM and SETM provide good estimates. In the left panel of Table 2, we see that the estimated bias, and RMSE are all small. Regarding the estimation efficiency, we observe that the SETM outperforms DETM, albeit marginally. The efficiency gain is due to that the SETM is a reduced model of DETM. For the DD data, the DETM method still performs well with RMSE shrinking toward 0 as the sample size increases. However, the SETM fails due to model misspecification as illustrated in the right panel of Table 2.
Table 2: Sample means and sample root mean square error (RMSE, in parenthesis) of $\pi$ estimates.
$n$ CD Data DD Data SETM DETM SETM DETM $\pi=0.3$ $\pi=0.7$ $\pi=0.3$ $\pi=0.7$ $\pi=0.3$ $\pi=0.7$ $\pi=0.3$ $\pi=0.7$ 1000 0.299 (0.016) 0.701(0.019) 0.300 (0.019) 0.703(0.021) 0.001(0.299) 0.001(0.699) 0.301(0.070) 0.672(0.078) 2000 0.300 (0.012) 0.701(0.013) 0.300 (0.013) 0.701(0.014) 0.001(0.299) 0.001(0.699) 0.300(0.042) 0.687(0.053) 3000 0.300 (0.009) 0.700(0.010) 0.300 (0.010) 0.700(0.012) 0.001(0.299) 0.001(0.699) 0.300(0.031) 0.687(0.044) 4000 0.300 (0.008) 0.700(0.009) 0.300 (0.009) 0.700(0.010) 0.001(0.299) 0.000(0.700) 0.299(0.026) 0.691(0.033) 5000 0.300 (0.007) 0.700(0.008) 0.300 (0.008) 0.700(0.009) 0.001(0.299) 0.000(0.700) 0.300(0.023) 0.694(0.028)
@@
We also construct two-sided $95\%$ confidence intervals for $\pi$, and the results are given in Figure 2. We report the estimated coverage probabilities in each of the different scenarios: 1) SETM on the CD PU data, 2) DETM on the CD PU data, 3) SETM on the DD PU data, and 4) DETM on the DD PU data. For the CD data, the coverage probabilities of both DETM and SETM reach the nominal level of 95% as the sample size increases. However, when the data are DD, the confidence intervals constructed using the SETM fail to cover the true parameter value $\pi$, resulting in a coverage probability that is close to 0. On the other hand, the DETM’s confidence intervals still perform well, remaining around the nominal levels of 95%.
Figure 2: Coverage probabilities of 95% confidence intervals for $\pi$.
@@
Lastly, we assess the prediction performance of our proposed method. We generate the CD and DD data the same way as in the “Estimations and Confidence Intervals for the Mixture Proportion” part, where $\pi\in\{0.3,0.7\}$ and $n=m\in\{1000,2000,3000,4000,5000\}$. To avoid overfitting, we generate a new validation dataset from the target distribution $p_{t}({\bf x},y)$ with sample size $n$, and the classifiers are applied to the validation dataset to calculate the classification accuracy.
We compare the following methods, including ours and some existing methods in the machine learning literature. 1) DETM: the proposed approximated Bayes classifier by replacing parameters with the MELEs under the DETM. 2) SETM: the approximated Bayes classifier by replacing parameters with the MELEs under the SETM. 3) PS: proposed by Bekker et al., (2019) and is based on the propensity score function. 4) KM2: proposed by Ramaswamy et al., (2016) using the reproducing kernel Hilbert space and is based on CD PU data. 5) TICE: proposed by Bekker and Davis, (2018) and is based on CD PU data.
Figure 3 illustrates the classification accuracies using 100 replicates. When the data are CD and $\pi=0.3$, all classifiers perform well, achieving an accuracy exceeding 95% in all sample sizes. However, when $\pi=0.7$, the KM2 method has the lowest accuracy at around 87%, while all other methods maintain at least 92% accuracy in all sample sizes.
@@
For the DD data, the simulation studies highlight discrepancies among different methods. In both $\pi=0.3$ and $\pi=0.7$ scenarios, our proposed DETM method achieves the highest classification rate of around 90%, with accuracy improving as the sample size increases. In comparison, at $\pi=0.3$, all benchmark methods have accuracies between 75% and 81%, with SETM performing the worst across all sample sizes. When $\pi=0.7$, the accuracies of benchmark methods PS, KM2, and TICE are around 60%. SETM again has the worst performance.
@@
The simulation study demonstrates that our proposed DETM method excels in estimation, inference, and classification tasks, regardless of whether DD or CD data are used. Under correct model assumptions, the SETM method performs well, with slightly improved estimation efficiency compared to DETM. However, SETM lacks robustness and performs poorly across all tasks when the CD assumption is violated. Additionally, the goodness-of-fit test effectively distinguishes DD data by accurately identifying CD data with a type-I error rate close to the nominal level and exhibiting high power in detecting non-CD data. For classification tasks, the proposed DETM method also shows superior performance compared to existing methods. We also conduct additional simulations for extreme scenarios, such as when $\pi$ is close to 0. However, due to space limitations, these results are included in the supplementary materials.
@@ 5.2 Mobile Phone Dataset
Before the analysis, we first preprocess the dataset to fit our needs. We first merge classes 0 and 1 to form the target positive data from $p_{t}({\bf x}|y=1)$ (low-end phone). We then use class 2 as the source positive data from $p_{s}({\bf x}|y=1)$ (low-end phone) and class 3 as target negative data from $p_{t}({\bf x}|y=0)$ (high-end phone). Then, the source positive data have a sample size of $n=500$, and the target data have a sample size of $m=1500$, where $1000$ are positive and $500$ are negative. Our formulation ensures that the low-end phones in the source and target are different, as low-end phones in the target data are less expensive; thus, the data can be seen as DD PU.
We first conduct a goodness-of-fit test for SETM vs. DETM. The value of the ELR statistic is $R_{N}=1302.222$, which gives a p-value of 0, calibrated by the limiting distribution $\chi^{2}_{20}$. Hence, we reject the null hypothesis with overwhelming evidence and adopt the DETM. The test results also align with the way we construct the PU data. Next, we examine the estimation and inference of the mixture proportion $\pi$. The true $\pi$ in the target data is $\pi^{o}=2/3$, and the estimate $\widehat{\pi}=0.6667$ is close to the true value. For the inference of $\pi$, we examine the behavior of the ELR function $R_{N}^{\ast}(\pi)$ in Figure 4. The vertical solid blue line is the true value of $\pi$. The red horizontal line is the 95% quantile of $\chi^{2}_{1}$, and the two dashed lines indicate the two endpoints of a $95\%$ confidence interval. We can see that the true value is within the confidence interval $[0.6425,0.6902]$.
@@
Lastly, we focus on the classification. Recall that the sample size of the positive target, positive source, and negative target data are $n_{1}=1000,n_{2}=500$, and $n_{3}=500$, respectively. To avoid overfitting, we reserve 20% of the positive target and negative target data and combine them for verification. Thus, the PU data have a sample size of $N=n+m$, where $n=n_{2}=500$ and $m=0.8n_{1}+0.8n_{3}=1200$. The sample size of the verification dataset is $0.2n_{1}+0.2n_{3}$. We repeated the data partition on the target data for 20 times and performed classification for each replicate. We compared the proposed DETM method with the benchmarks used in the simulation study. Table 3 summarizes the classification accuracy averaged from $20$ sample splitting replicates. The numbers in the parenthesis are the standard deviation of the 20 accuracies from 20 replicates. Our proposed method, DETM, has surprisingly achieved perfect classification, while the benchmark methods have accuracies of about 88. 20% to 91. 30%.
@@
Method DETM PS KM2 TICE Accuracy 100% (0%) 91.30% (1.39%) 89.05% (3.31%) 88.20% (1.46%)
@@ 6 Concluding Remarks
We would like to compare our transfer learning framework with the propensity score framework used in the literature (Bekker et al., 2019; Bekker and Davis, 2020) in this final section. Unlike the transfer learning setting, which involves two samples, the propensity score framework considers only a single sample, where both labeled and unlabeled data are assumed to be generated from a joint distribution $p({\bf x},y)$. In this framework, an additional binary random variable $S$ is introduced to determine whether an example is labeled ($S=1$) or unlabeled ($S=0$). The conditional distribution for positive examples, $\Pr(S=1|{\bf x},y=1)$, is called the propensity score function. It is unnecessary to consider $y=0$, as $\Pr(S=1|{\bf x},y=0)=0$. If the propensity score function does not depend on ${\bf x}$ (i.e., $\Pr(S=1|{\bf x},y=1)$ is a constant), we say the PU data are selected completely at random (SCAR). Otherwise, the PU data are selected at random (SAR).
The SCAR and SAR PU settings correspond to the CD and DD assumptions, respectively. To illustrate this, we can write the propensity score function as
$\begin{split}&\Pr(S=1|{\bf x},y=1)=\frac{p(S=1,{\bf x},y=1)}{p({\bf x},y=1)}\\ =&\frac{p({\bf x}|y=1,S=1)\Pr(S=1,y=1)}{p({\bf x}|y=1,S=0)\Pr(S=0,y=1)+p({\bf x}|y=1,S=1)\Pr(S=1,y=1)}\\ =&\frac{\Pr(S=1,y=1)}{\dfrac{p({\bf x}|y=1,S=0)}{p({\bf x}|y=1,S=1)}\Pr(S=0,y=1)+\Pr(S=1,y=1)}.\end{split}$ (13)
Notice that $p_{s}({\bf x}|y=1)=p({\bf x}|y=1,S=1)$ and $p_{t}({\bf x}|y=1)=p({\bf x}|y=1,S=0)$. Then from (13) we can see that the CD assumption is equivalent to the SCAR assumption. Also, when the propensity score does depend on ${\bf x}$ (i.e., SAR PU), it implies that the CD assumption does not hold. In other words, the SAR PU setting implies the DD PU setting under the transfer learning framework.
What does the DETM imply about the propensity score function? From (13), we can see that under the DETM, the propensity score function follows a logistic regression model:
@@
$\Pr(S=1|{\bf x},y=1)=\frac{\exp(\alpha^{\prime}_{1}+{\bf x}^{\rm T}{\boldsymbol{\beta}}_{1})}{1+\exp(\alpha^{\prime}_{1}+{\bf x}^{\rm T}{\boldsymbol{\beta}}_{1})}~{}~{}\mbox{with}~{}~{}\alpha^{\prime}_{1}=\alpha_{1}+{\rm log}\left\{\frac{\Pr(S=1,y=1)}{\Pr(S=0,y=1)}\right\}.$
We conclude this paper by highlighting several promising directions for future research. One intriguing extension involves exploring the open set label shift setting (e.g., Garg et al., 2022), where the target domain introduces a new class not present in the source domain, but the source domain has multiple classes. This setting expands the scope of PU data and presents new challenges and opportunities for DETM to demonstrate its adaptability and effectiveness. Another critical area for future investigation is the robustness of DETM in the presence of noisy labels. Enhancing the model’s resilience to such imperfections is vital for ensuring reliable performance in real-world applications, where data quality can often be compromised. Furthermore, if the dimension of ${\bf X}$ is high, penalized methods may be considered before fitting the ETMs. Through the estimation procedures presented in Section 3.2, particularly Step 2 in the M-step, we may, for instance, use glmnet to easily incorporate penalized maximum (empirical likelihood) methods to deal with high-dimensional ${\bf X}$.
Supplementary Materials
The online supplementary document includes all technical details and proofs, additional details on the M-step of the EM algorithm, form of ${\boldsymbol{\Sigma}}$ and conditions, extra simulation studies, and features of the mobile phone dataset. We also provide the R package and all necessary code to reproduce the analysis.
Acknowledgment
We sincerely thank the joint editor and associate editor for their time and effort in overseeing the review of this manuscript. We are also grateful to the three anonymous reviewers for their insightful comments, which greatly improved our work. Additionally, we appreciate Dr. Xiyue Han for his helpful discussions in the early stages of this research. Finally, we extend our gratitude to Professor Yukun Liu for his valuable and constructive feedback. This work is partially supported by the Natural Sciences and Engineering Research Council of Canada (RGPIN-2023-03479 and RGPIN-2020-04964) and the Faculty of Mathematics Research Chair Funding of the University of Waterloo.
The majority of the work was completed during Siyan Liu’s visit to the University of Waterloo as a visiting PhD student under the joint supervision of Qinglong Tian and Pengfei Li. Qinglong Tian is the corresponding author.
Disclosure Statement
No potential conflict of interest was reported by the authors.